Shift report 20000811 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Fri Aug 11 2000 - 16:44:33 EDT

  • Next message: Hans Boggild, NBI, Copenhagen: "brahms runplan as of 11/8"

    ------------------------------------------------------------
    Date: Fri Aug 11 16:44:33 2000, Shift: 08:00-16:00
    Supervisor: RD, with: ZIS
    Log book#4 pages 49 - 55
    
    Shift Summary:
    ------------------------------------------------------------
    1. DEAD
    2. run 2188 ZDC LED before work
       run 2196 ZDC LED after work
    3. Today we looked at LED response in ZDC to identify a 
    possible gain shift. The result of the study is that there is no shift.
    The single neutron peak became more difficult to see in the
    ntuple plots because of additional background that must be related
    to tune of the machine.
    During this work we checked DC offfsets of sum circuits, and
    made the Camac ADC gate for the unattenuated signal 10 nsec
    earlier.
    
    The Forward spectrometer has been moved from 8 to 5 degrees.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Fri Aug 11 2000 - 16:47:58 EDT